# TNO, The Hague, The Netherlands

### v1

@@ 3.2.1 VRP and TSP quantum solutions
The work by Spyridis et al. spyridis2023variational , focuses on the Multiple Travelling Salesman Problem (m-TSP). The study explores the use of the QAOA to efficiently handle the m-TSP, formulating a QUBO problem. The authors highlight potential applications in IoT systems, emphasising the need for innovative error correction techniques to address limitations of NISQ devices. An other variant of TSP is the selective travelling salesman problem (sTSP), which is also known as the orienteering problem (OP), where each destination is associated with a prize. Given a budget that must not be exceeded, the salesperson is not necessarily required to visit all the cities, and his mission is to collect as many prizes as possible. In Le et al. le2023quantum , the authors provide a QUBO formulation for this problem and perform experimental results include an analysis of parameters such as annealing time, chain strength, and solution quality for different instances of sTSP with varying numbers of cities.
@@
Salehi et al. salehi2022unconstrained and Papalitsas et al. papalitsas2019qubo both look at the Travelling Salesman Problem with Time Windows (TSPTW). The TSPTW involves finding a tour with the minimum cost, where each city must be visited within a specified time window. Papalitsas et al. give a first QUBO formulation of this problem. Salehi et al. introduce three different formulations. The first formulation considers both earliest start times and due times for each city. The second formulation presents a higher-order binary model that is more space-efficient. The third formulation is based on an alternative QUBO formulation derived from integer linear programming. These formulations allow for multiple assignments of binary variables to encode the optimal route without penalty, and they can be adapted for other variants of the TSP. The study by Salehi et al. investigates the efficiency of edge-based and integer linear programming (ILP) formulations through experiments on the D-Wave Advantage quantum annealer. The results indicate that the ILP model outperforms the edge-based model, particularly for instances with four cities, showcasing a higher probability of obtaining samples encoding the optimal route. The paper suggests a natural progression for future work, involving a more nuanced exploration of penalty values for time window constraints and the investigation of alternative techniques for quadratisation in the node-based formulation.
@@
Marsoit et al. marsoit2021quantum address the VRP with uncertain data, proposing a quantum computing approach. In real-world scenarios there are uncertainties, which impact the effectiveness of VRP solutions, think of varying travel times due to traffic congestion, uncertain customer demands, service disruptions, or weather conditions. Failing to account for these uncertainties may lead to inefficient route planning, missed time windows, increased operational costs, and customer dissatisfaction. The study formulates the problem, compares it with classical optimisation challenges, and claims to demonstrate the effectiveness of the quantum approach through a numerical example. However, it is not clear which quantum approach they use. The research acknowledges current limitations of quantum hardware and calls for further advancements.
@@
Dixit et al. dixit2023quantum2 also discuss VRP with uncertainty, addressing the Stochastic Time Dependent Shortest Path (STDSP) problem. In STDSP, uncertainties arise from factors like demand and supply fluctuations, making the representation of link travel time as a random variable necessary. The study introduces a Quadratic Constrained Binary Optimisation Problem for the STDSP. The research explores the efficiency of QA in comparison to classical solvers, specifically CPLEX. The results indicate that quantum computing, despite its current limitations in terms of qubit numbers and noise susceptibility, exhibits linear computational efficiency.
@@
Fitzek et al. fitzek2021applying explore the use of QAOA for the Heterogeneous Vehicle Routing Problem (HVRP). The HVRP involves determining optimal routes for a fleet of vehicles with varying capacities to deliver goods to customers. They formulate the HVRP as an Ising Hamiltonian, showcasing scalability with increasing problem size. The study employs QAOA, which requires defining a cost function and proposing an ansatz for optimization. The study emphasises the trade-off between classical optimiser performance and runtime, providing insights into the potential of quantum computing for solving large instances of HVRP with advanced hardware.
@@
Lo et al. lo2021genetic introduce a quantum random number generator (QRNG) in solving pollution-routing problems (PRPs) for sustainable logistics. The hybrid model incorporates a modified k-means algorithm and a genetic algorithm, showcasing the applicability of quantum random number generation in optimising routes for emission reduction. The study suggests future research directions involving metaheuristic algorithms and advancements in quantum computing technology.
@@
The work by Harikrishnakumar harikrishnakumar2020quantum introduces a QA approach to address the Multi-Depot Capacitated Vehicle Routing Problem (MDCVRP). The MDCVRP involves assigning routes to vehicles from multiple depots, each with varying capacities, to serve spatially distributed customer locations while satisfying capacity constraints. The authors formulate the MDCVRP and its dynamic version, the D-MDCVRP, as QUBO problems. The authors discuss the complexity of the formulated problems, considering the total number of decision variables. They also present a step-by-step solution framework for solving QUBO formulations on quantum annealing hardware. The authors outline future work, which includes solving the QUBO formulations on D-Wave (here the 2000Q), comparing results against classical heuristic algorithms like Tabu Search, and investigating the scalability of the quantum annealing approach concerning the number of vehicles and depots.
@@
The works of Warren warren2020solving and Osaba et al.osaba2022systematic gives an overview of VRP and TSP implementations. Warren conducts an analysis of four software programs designed to solve the symmetric TSP, among which D-Wave’s solutions and the work by Feld feld2019hybrid we saw before. Osaba et al give three main observations in the context of quantum optimisation and routing problems. Regarding the current state of routing problems they indicate that traditional routing problems like the TSP and VRP have been extensively studied, but their formulations often fall short of addressing advanced or realistic scenarios. Researchers face challenges in tailoring problems to hardware capacities, leading to limitations in addressing complex formulations. The second observation is the current state of the hardware. current commercial quantum devices have inherent limitations such as noise and decoherence. Researchers have worked on error correction and mitigation strategies, but the focus on this aspect is relatively low. The last observation is on the parameterisation sensitivity in quantum devices. Quantum devices are highly sensitive to parameterisation in various aspects, including problem formulation, algorithm tuning, and hardware specifications. Despite this sensitivity, a majority of practical papers focus on solving specific problems without deep exploration of parameter fine-tuning. Proper parameterisation, especially the choice of (QUBO) penalty values in problem formulation, is highlighted as a significant challenge.
@@
Last, the work by Poggel et al. poggel2023recommending introduces a framework designed to automate and optimise the decision-making process in quantum-assisted optimisation for real-world problems. Addressing the challenges of formulating, encoding, and selecting algorithms and hardware options, the proposed abstraction layer aims to make quantum-computing-assisted solution techniques accessible to end-users without requiring in-depth knowledge of quantum technologies. The framework incorporates state-of-the-art hybrid algorithms, encoding, and decomposition techniques in a modular manner, allowing evaluation using problem-specific performance metrics. Tools for graphical analysis of variational quantum algorithms are developed, and classical, fault-tolerant quantum, and quantum-inspired methods can be included for a fair comparison. The approach is demonstrated and validated on the CVRP.
@@ 3.2.2 Other Routing problems
In addition to addressing the classic Vehicle Routing Problem (VRP) and Travelling Salesman Problem (TSP), quantum computing has been applied to various other routing problems, showcasing its potential in optimising complex logistical scenarios. The first example is the shipment rerouting problem by Yarkoni et al. yarkoni2021solving , addressing the challenge of partially filled trucks transporting shipments between a network of hubs. The goal is to minimise the total distance travelled by these trucks, thus reducing fuel consumption and enhancing cost efficiency. The problem instances and optimisation techniques are based on real-world data from an existing shipment network in Europe. The authors create a QUBO formulation and then employ both classical and hybrid quantum-classical algorithms to solve these QUBOs. The quantum platforms, particularly D-Wave, are found to be promising for addressing this logistics problem due to their qubit support, compared to gate-based and quantum-inspired optimisers. The article mentions that due to qubit limitations in current quantum hardware, the experiments are conducted with minimal parameters. It anticipates that as quantum hardware improves, more complex real-world problems can be solved using quantum computing.
@@
Azzaoui et al. azzaoui2021quantum extended the scope to smart logistics systems, proposing Quantum Approximate Optimization Algorithm (QAOA) for route optimisation. They emphasise the importance of considering, next to cost and time, also carbon footprint in logistics operations. They propose the deployment of QAOA to calculate and optimise travel paths, selecting the best route to reduce costs, time and carbon footprint, in a multi objective approach, in smart logistics systems. Next to this approach, they suggest implementing a private blockchain network in smart logistics systems to ensure secure and private communication between different tiers of the smart logistics system, creating a trustworthy cluster. They then conduct simulations, using IBM Quantum Lab and Qiskit, along with blockchain-based secure communication simulations using IBM Hyperledger. The results show promising outcomes, including a 91% success rate in route optimisation, a 10% cost reduction, a 6% reduction in carbon footprint, and improved system scalability. The paper concludes by highlighting the potential benefits of the proposed framework, aiming to improve the scalability, reduce the carbon footprint, and cut down transportation and logistics costs in smart logistics systems.
@@
The article by Atchade et al. atchade2021qrobot introduces quantum computing to the field of robotics, specifically focusing on optimising distance travelled by the robots in warehouses and distribution centres where order picking and batching is performed. The authors develop a real-time quantum optimisation algorithm for this purpose, implemented on a Raspberry Pi 4 as a proof of concept. The proposed system aims to minimise distance travelled and optimise order batching. The system allows for hybrid computing, combining quantum and classical processing. The environment developed enables the execution of quantum algorithms on IBMQ, Amazon Braket (D-Wave), and Pennylane, either locally or remotely. The study indicates that Amazon Braket has better time performance than Qiskit or Pennylane. They discuss potential improvements and considerations for applying the system to scenarios with multiple robots and batches.
@@
In the aviation sector, Makhanov et al. makhanov2023quantum dive into the complexities of optimising flight paths, emphasising the use of quantum computing to overcome computational challenges associated with this crucial aerospace engineering operation. The research introduces a customisable modular framework designed to accommodate specific simulation requirements. It also examines the application of quantum-enhanced algorithms using Grover’s algorithm, on various quantum architectures and simulations using CPUs and GPUs. The study presents results from running quantum algorithms on IBM hardware, highlighting the potential advantages and challenges associated with quantum computing in the aerospace industry and acknowledging the need for further innovation to practically achieve the theoretical speedup promised by quantum algorithms.
@@
The paper by Neukart et al. neukart2017traffic focuses on a real-world application, namely traffic flow optimisation in dense road networks. The authors describe the process of mapping a traffic flow optimisation problem onto a QUBO formulation, to demonstrate the feasibility of using quantum computing for real-time, time-critical optimisation tasks, such as continuous redistribution of position data for cars in busy road networks. Due to the limited size and connectivity of current D-Wave’s QPUs, a hybrid quantum and classical approach is employed to address the traffic flow problem. The quality of solutions is assessed by counting the number of congested roads after optimisation. The study compares the outcomes against random assignments of routes. The results indicate that the hybrid quantum approach redistributes traffic in a way that reduces the number of congested roads compared to both random assignments and the original route assignment.
@@
Last example in the routing section, is the work by Phillipson et al. phillipson2021multimodal . Also this study presents a real-world application in multimodal container planning and demonstrates how to map this problem to a QUBO formulation for practical implementation on the quantum annealer produced by D-Wave. It describes the implementation of the formulation on both SA and D-Wave’s QPU. Due to current limitations in the size of quantum computers, the authors initially formulate the problem with a limited number of decision variables and later propose expansions for larger-scale applications. The conclusion emphasises the need for re-formulating real-world problems to match the QUBO formulation and highlights the potential of quantum computing for solving such problems as the technology matures. The implementation issues on D-Wave’s QA, such as finding the right embedding, defining chain strength, and penalty functions, are addressed through a parameter grid search. The paper recommends further research on finding more general methods for these parameters, extending the set with alternative paths, and introducing a quantum variant of column generation for optimising paths.
@@ 3.3 Logistic Network Design
At the heart of logistical optimisation lies the strategic design of logistic networks, a complex process that involves careful planning, analysis, and decision-making. Logistic Network Design (LND) explores how organisations can craft and refine their supply chain infrastructure to enhance efficiency, reduce costs, and meet the ever-evolving demands of the market. From the selection of distribution centres to the configuration of transportation routes, this exploration aims to unravel the strategic considerations that underpin a resilient and responsive logistical framework. One important classical problem within this area is the Facility Location Problem. For this, Mahasinghe mahasinghe2021qubo gives a basic QUBO formulation.
@@
The first work on LND by Ding et al. ding2021implementation ; ding2019logistic discusses the application of quantum annealing to address logistic network design problems (NDPs), which are described as abstract optimisation problems aimed at finding the optimal configuration of supply chain infrastructures and facilities based on customer demand, while minimising costs. They translate the cost function with constraints into an Ising problem and benchmark their results by measuring the accuracy of the solutions against optimal published solutions, with an average error of less than 1%. Additionally, they compare the performance of their quantum approach with classical algorithms and observe a significant reduction in the number of iterations required. The article emphasises that, even though current quantum annealers are not yet achieving quantum supremacy, they can effectively address relevant supply-chain problems. Following the authors, this demonstrates the potential for state-of-the-art quantum annealers to encode and solve complex logistics problems, despite limitations in quantum hardware. Furthermore, the authors introduce an alternative approach called the combined quantum annealing algorithm, which utilises two layers with feedback-control interaction between them. This approach is tested on 12 NDP instances using both quantum annealer simulators and the D-Wave QA, yielding positive results in terms of accuracy and performance compared to classical algorithms.
@@
The work by Dixit et al. dixit2023quantum discusses the Transport Network Design Problem (TNDP), involving optimising network capacities, scheduling maintenance, and identifying new links under resource constraints. Traditionally, meta-heuristic methods like Tabu Search are employed. The paper formulates TNDP as a bi-level problem, with the upper level represented as a QUBO problem, solved using quantum annealing on D-Wave’s QA. The results are compared with Tabu Search, demonstrating significant computational benefits using quantum annealing.
@@
Next, Malviya et al. malviya2023logistics explore the potential of quantum annealing for a parcel distribution centre network optimisation using data instances of different sizes. Using a small data instance that fits the current technology, they use D-Wave’s built-in hybrid quantum-classical Kerberos sampler via AWS Braket for solving the problem. An iterative hybrid approach for improving the quality of solutions obtained from the Kerberos sampler is used and then the solutions are compared to the random sampler, a greedy heuristic, an exact solver and SA. Improvement in objective value over random has been shown for different approaches over different data instance sizes. Kerberos sampler performs better than greedy heuristic and at par with exact solver for some data instances.
@@
Gabbassov et al. gabbassov2022transit address the challenge of balancing service efficiency and accessibility in urban transit facility planning. The focus is on transit facility consolidation as a cost-effective strategy to enhance service quality. For this, they propose an optimisation framework that integrates Geographical Information Systems (GIS), decision-making analysis, and quantum technologies. The framework includes a mathematical model capturing non-linear interactions between facilities, demand nodes, inter-facility competition, ridership demand, and spatial coverage, solving the Spatial Interaction Coverage (SIC) problem. The QUBO based model can, as they suggest, be implemented on various quantum optimisation metaheuristics, including QA, quantum-inspired Digital Annealing, Coherent Ising Machines, and universal GBC. The framework is applied to the public transit facility redundancy problem in the British Columbia Vancouver metropolitan area and demonstrates a 40% reduction in the number of facilities while maintaining the same service accessibility, thereby improving efficiency, using a D-Wave Hybrid quantum-classical solver. Additionally, numerical experiments on a synthetically generated dataset demonstrate that the Hybrid solver yields statistically superior demand coverage compared to a SIC problem solved with the APOPT [^2] Advanced Process OPTimizer solver.
@@
The paper by Klar et al. klar2022quantum addresses the challenges of factory layouts, emphasising their impact on operational costs. The problem involves assigning functional units to positions, optimising transportation distances. Manual layout planning is time-consuming due to the complexity of considering multiple parameters simultaneously. The study introduces a QA-based approach for factory layout planning, involving the formulation of the problem as a QUBO problem. The results show that QA can solve layout planning problems of different sizes within seconds, overcoming the trade-off between solution quality and computation time.

### v2

@@ 3.2.1 VRP and TSP quantum solutions
The work by Spyridis et al. [146], focuses on the Multiple Travelling Salesman Problem (m-TSP). The study explores the use of the QAOA to efficiently handle the m-TSP, formulating a QUBO problem. The authors highlight potential applications in IoT systems, emphasising the need for innovative error correction techniques to address limitations of NISQ devices. An other variant of TSP is the selective travelling salesman problem (sTSP), which is also known as the orienteering problem (OP), where each destination is associated with a prize. Given a budget that must not be exceeded, the salesperson is not necessarily required to visit all the cities, and his mission is to collect as many prizes as possible. In Le et al. [81], the authors provide a QUBO formulation for this problem and perform experimental results include an analysis of parameters such as annealing time, chain strength, and solution quality for different instances of sTSP with varying numbers of cities.
@@
Salehi et al. [128] and Papalitsas et al. [110] both look at the Travelling Salesman Problem with Time Windows (TSPTW). The TSPTW involves finding a tour with the minimum cost, where each city must be visited within a specified time window. Papalitsas et al. give a first QUBO formulation of this problem. Salehi et al. introduce three different formulations. The first formulation considers both earliest start times and due times for each city. The second formulation presents a higher-order binary model that is more space-efficient. The third formulation is based on an alternative QUBO formulation derived from integer linear programming. These formulations allow for multiple assignments of binary variables to encode the optimal route without penalty, and they can be adapted for other variants of the TSP. The study by Salehi et al. investigates the efficiency of edge-based and integer linear programming (ILP) formulations through experiments on the D-Wave Advantage quantum annealer. The results indicate that the ILP model outperforms the edge-based model, particularly for instances with four cities, showcasing a higher probability of obtaining samples encoding the optimal route. The paper suggests a natural progression for future work, involving a more nuanced exploration of penalty values for time window constraints and the investigation of alternative techniques for quadratisation in the node-based formulation.
@@
Marsoit et al. [92] address the VRP with uncertain data, proposing a quantum computing approach. In real-world scenarios there are uncertainties, which impact the effectiveness of VRP solutions, think of varying travel times due to traffic congestion, uncertain customer demands, service disruptions, or weather conditions. Failing to account for these uncertainties may lead to inefficient route planning, missed time windows, increased operational costs, and customer dissatisfaction. The study formulates the problem, compares it with classical optimisation challenges, and claims to demonstrate the effectiveness of the quantum approach through a numerical example. However, it is not clear which quantum approach they use. The research acknowledges current limitations of quantum hardware and calls for further advancements.
@@
Dixit et al. [46] also discuss VRP with uncertainty, addressing the Stochastic Time Dependent Shortest Path (STDSP) problem. In STDSP, uncertainties arise from factors like demand and supply fluctuations, making the representation of link travel time as a random variable necessary. The study introduces a Quadratic Constrained Binary Optimisation Problem for the STDSP. The research explores the efficiency of QA in comparison to classical solvers, specifically CPLEX. The results indicate that quantum computing, despite its current limitations in terms of qubit numbers and noise susceptibility, exhibits linear computational efficiency.
@@
Fitzek et al. [50] explore the use of QAOA for the Heterogeneous Vehicle Routing Problem (HVRP). The HVRP involves determining optimal routes for a fleet of vehicles with varying capacities to deliver goods to customers. They formulate the HVRP as an Ising Hamiltonian, showcasing scalability with increasing problem size. The study employs QAOA, which requires defining a cost function and proposing an ansatz for optimization. The study emphasises the trade-off between classical optimiser performance and runtime, providing insights into the potential of quantum computing for solving large instances of HVRP with advanced hardware.
@@
Lo et al. [85] introduce a quantum random number generator (QRNG) in solving pollution-routing problems (PRPs) for sustainable logistics. The hybrid model incorporates a modified k-means algorithm and a genetic algorithm, showcasing the applicability of quantum random number generation in optimising routes for emission reduction. The study suggests future research directions involving metaheuristic algorithms and advancements in quantum computing technology.
@@
The work by Harikrishnakumar [62] introduces a QA approach to address the Multi-Depot Capacitated Vehicle Routing Problem (MDCVRP). The MDCVRP involves assigning routes to vehicles from multiple depots, each with varying capacities, to serve spatially distributed customer locations while satisfying capacity constraints. The authors formulate the MDCVRP and its dynamic version, the D-MDCVRP, as QUBO problems. The authors discuss the complexity of the formulated problems, considering the total number of decision variables. They also present a step-by-step solution framework for solving QUBO formulations on QA hardware. The authors outline future work, which includes solving the QUBO formulations on D-Wave (here the 2000Q), comparing results against classical heuristic algorithms like Tabu Search, and investigating the scalability of the QA approach concerning the number of vehicles and depots.
@@
The works of Warren [158] and Osaba et al.[107] gives an overview of VRP and TSP implementations. Warren conducts an analysis of four software programs designed to solve the symmetric TSP, among which D-Wave’s solutions and the work by Feld [49] we saw before. Osaba et al give three main observations in the context of quantum optimisation and routing problems. Regarding the current state of routing problems they indicate that traditional routing problems like the TSP and VRP have been extensively studied, but their formulations often fall short of addressing advanced or realistic scenarios. Researchers face challenges in tailoring problems to hardware capacities, leading to limitations in addressing complex formulations. The second observation is the current state of the hardware. current commercial quantum devices have inherent limitations such as noise and decoherence. Researchers have worked on error correction and mitigation strategies, but the focus on this aspect is relatively low. The last observation is on the parameterisation sensitivity in quantum devices. Quantum devices are highly sensitive to parameterisation in various aspects, including problem formulation, algorithm tuning, and hardware specifications. Despite this sensitivity, a majority of practical papers focus on solving specific problems without deep exploration of parameter fine-tuning. Proper parameterisation, especially the choice of (QUBO) penalty values in problem formulation, is highlighted as a significant challenge.
@@
Lastly, the work by Poggel et al. [117] introduces a framework designed to automate and optimise the decision-making process in quantum-assisted optimisation for real-world problems. Addressing the challenges of formulating, encoding, and selecting algorithms and hardware options, the proposed abstraction layer aims to make quantum-computing-assisted solution techniques accessible to end-users without requiring in-depth knowledge of quantum technologies. The framework incorporates state-of-the-art hybrid algorithms, encoding, and decomposition techniques in a modular manner, allowing evaluation using problem-specific performance metrics. Tools for graphical analysis of variational quantum algorithms are developed, and classical, fault-tolerant quantum, and quantum-inspired methods can be included for a fair comparison. The approach is demonstrated and validated on the CVRP.
@@ 3.2.2 Other Routing problems
In addition to addressing the classic Vehicle Routing Problem (VRP) and Travelling Salesman Problem (TSP), quantum computing has been applied to various other routing problems, showcasing its potential in optimising complex logistical scenarios. The first example is the shipment re-routing problem by Yarkoni et al. [168], addressing the challenge of partially filled trucks transporting shipments between a network of hubs. The goal is to minimise the total distance travelled by these trucks, thus reducing fuel consumption and enhancing cost efficiency. The problem instances and optimisation techniques are based on real-world data from an existing shipment network in Europe. The authors create a QUBO formulation and then employ both classical and hybrid quantum-classical algorithms to solve these QUBOs. The quantum platforms, particularly D-Wave, are found to be promising for addressing this logistics problem due to their qubit support, compared to gate-based and quantum-inspired optimisers. The article mentions that due to qubit limitations in current quantum hardware, the experiments are conducted with minimal parameters. It anticipates that as quantum hardware improves, more complex real-world problems can be solved using quantum computing.
@@
Azzaoui et al. [13] extended the scope to smart logistics systems, proposing Quantum Approximate Optimization Algorithm (QAOA) for route optimisation. They emphasise the importance of considering, next to cost and time, also carbon footprint in logistics operations. They propose the deployment of QAOA to calculate and optimise travel paths, selecting the best route to reduce costs, time and carbon footprint, in a multi objective approach, in smart logistics systems. Next to this approach, they suggest implementing a private blockchain network in smart logistics systems to ensure secure and private communication between different tiers of the smart logistics system, creating a trustworthy cluster. They then conduct simulations, using IBM Quantum Lab and Qiskit, along with blockchain-based secure communication simulations using IBM Hyperledger. The results show promising outcomes, including a 91% success rate in route optimisation, a 10% cost reduction, a 6% reduction in carbon footprint, and improved system scalability. The paper concludes by highlighting the potential benefits of the proposed framework, aiming to improve the scalability, reduce the carbon footprint, and cut down transportation and logistics costs in smart logistics systems.
@@
The article by Atchade et al. [11] introduces quantum computing to the field of robotics, specifically focusing on optimising distance travelled by the robots in warehouses and distribution centres where order picking and batching is performed. The authors develop a real-time quantum optimisation algorithm for this purpose, implemented on a Raspberry Pi 4 as a proof of concept. The proposed system aims to minimise distance travelled and optimise order batching. The system allows for hybrid computing, combining quantum and classical processing. The environment developed enables the execution of quantum algorithms on IBMQ, Amazon Braket (D-Wave), and Pennylane, either locally or remotely. The study indicates that Amazon Braket has better time performance than Qiskit or Pennylane. They discuss potential improvements and considerations for applying the system to scenarios with multiple robots and batches.
@@
In the aviation sector, Makhanov et al. [87, 88] dive into the complexities of optimising flight paths, emphasising the use of quantum computing to overcome computational challenges associated with this crucial aerospace engineering operation. The research introduces a customisable modular framework designed to accommodate specific simulation requirements. It also examines the application of quantum-enhanced algorithms using Grover’s algorithm, on various quantum architectures and simulations using CPUs and GPUs. The study presents results from running quantum algorithms on IBM hardware, highlighting the potential advantages and challenges associated with quantum computing in the aerospace industry and acknowledging the need for further innovation to practically achieve the theoretical speedup promised by quantum algorithms. Khan et al. [74] optimise flight routes using a combination of Grover’s Algorithm and the Quantum Approximate Optimization Algorithm (QAOA). Grover’s Algorithm speeds up the search for optimal routes, while QAOA refines these routes based on real-time data like weather and air traffic. This hybrid approach enhances fuel efficiency, operational effectiveness, and adaptability to dynamic conditions.
@@
The paper by Neukart et al. [104] focuses on a real-world application, namely traffic flow optimisation in dense road networks. The authors describe the process of mapping a traffic flow optimisation problem onto a QUBO formulation, to demonstrate the feasibility of using quantum computing for real-time, time-critical optimisation tasks, such as continuous redistribution of position data for cars in busy road networks. Due to the limited size and connectivity of current D-Wave’s QPUs, a hybrid quantum and classical approach is employed to address the traffic flow problem. The quality of solutions is assessed by counting the number of congested roads after optimisation. The study compares the outcomes against random assignments of routes. The results indicate that the hybrid quantum approach redistributes traffic in a way that reduces the number of congested roads compared to both random assignments and the original route assignment. This work continued by Salloum et al. [130] proposing a method to decompose large traffic problems into smaller sub-problems, leveraging QA for improved speed and effectiveness in managing urban traffic flow, demonstrated through experiments with up to 500 cars. The same authors explore in [131] optimizing urban traffic flow using Quantum Annealing (QA) to address congestion. Now they employ a hybrid quantum-classical approach, transitioning to a purely quantum method for smaller Quadratic Unconstrained Binary Optimization (QUBO) problems.
@@
Last example in the routing section, is the work by Phillipson et al. [113]. Also this study presents a real-world application in multimodal container planning and demonstrates how to map this problem to a QUBO formulation for practical implementation on the quantum annealer produced by D-Wave. It describes the implementation of the formulation on both SA and D-Wave’s QPU. Due to current limitations in the size of quantum computers, the authors initially formulate the problem with a limited number of decision variables and later propose expansions for larger-scale applications. The conclusion emphasises the need for re-formulating real-world problems to match the QUBO formulation and highlights the potential of quantum computing for solving such problems as the technology matures. The implementation issues on D-Wave’s QA, such as finding the right embedding, defining chain strength, and penalty functions, are addressed through a parameter grid search. The paper recommends further research on finding more general methods for these parameters, extending the set with alternative paths, and introducing a quantum variant of column generation for optimising paths.
@@ 3.3 Logistic Network Design
At the heart of logistical optimisation lies the strategic design of logistic networks, a complex process that involves careful planning, analysis, and decision-making. Logistic Network Design (LND) explores how organisations can craft and refine their supply chain infrastructure to enhance efficiency, reduce costs, and meet the ever-evolving demands of the market. From the selection of distribution centres to the configuration of transportation routes, this exploration aims to unravel the strategic considerations that underpin a resilient and responsive logistical framework. One important classical problem within this area is the Facility Location Problem. For this, Mahasinghe [86] gives a basic QUBO formulation.
@@
The first work on LND by Ding et al. [44, 43] discusses the application of QA to address logistic network design problems (NDPs), which are described as abstract optimisation problems aimed at finding the optimal configuration of supply chain infrastructures and facilities based on customer demand, while minimising costs. They translate the cost function with constraints into an Ising problem and benchmark their results by measuring the accuracy of the solutions against optimal published solutions, with an average error of less than 1%. Additionally, they compare the performance of their quantum approach with classical algorithms and observe a significant reduction in the number of iterations required. The article emphasises that, even though current quantum annealers are not yet achieving quantum supremacy, they can effectively address relevant supply-chain problems. Following the authors, this demonstrates the potential for state-of-the-art quantum annealers to encode and solve complex logistics problems, despite limitations in quantum hardware. Furthermore, the authors introduce an alternative approach called the combined QA algorithm, which utilises two layers with feedback-control interaction between them. This approach is tested on 12 NDP instances using both quantum annealer simulators and the D-Wave QA, yielding positive results in terms of accuracy and performance compared to classical algorithms.
@@
The work by Dixit et al. [45] discusses the Transport Network Design Problem (TNDP), involving optimising network capacities, scheduling maintenance, and identifying new links under resource constraints. Traditionally, meta-heuristic methods like Tabu Search are employed. The paper formulates TNDP as a bi-level problem, with the upper level represented as a QUBO problem, solved using QA on D-Wave’s quantum annealer. The results are compared with Tabu Search, demonstrating significant computational benefits using QA.
@@
Next, Malviya et al. [90] explore the potential of QA for a parcel distribution centre network optimisation using data instances of different sizes. Using a small data instance that fits the current technology, they use D-Wave’s built-in hybrid quantum-classical Kerberos sampler via AWS Braket for solving the problem. An iterative hybrid approach for improving the quality of solutions obtained from the Kerberos sampler is used and then the solutions are compared to the random sampler, a greedy heuristic, an exact solver and SA. Improvement in objective value over random has been shown for different approaches over different data instance sizes. Kerberos sampler performs better than greedy heuristic and at par with exact solver for some data instances. The same authors [89] presents a quantum approach using QAOA to redesign last mile delivery networks. It formulates the problem with hard constraints, optimizing distribution center operations and customer allocations. Results show improved efficiency and scalability for various problem instances.
@@
Gabbassov et al. [52] address the challenge of balancing service efficiency and accessibility in urban transit facility planning. The focus is on transit facility consolidation as a cost-effective strategy to enhance service quality. For this, they propose an optimisation framework that integrates Geographical Information Systems (GIS), decision-making analysis, and quantum technologies. The framework includes a mathematical model capturing non-linear interactions between facilities, demand nodes, inter-facility competition, ridership demand, and spatial coverage, solving the Spatial Interaction Coverage (SIC) problem. The QUBO based model can, as they suggest, be implemented on various quantum optimisation metaheuristics, including QA, quantum-inspired Digital Annealing, Coherent Ising Machines, and universal GBC. The framework is applied to the public transit facility redundancy problem in the British Columbia Vancouver metropolitan area and demonstrates a 40% reduction in the number of facilities while maintaining the same service accessibility, thereby improving efficiency, using a D-Wave Hybrid quantum-classical solver. Additionally, numerical experiments on a synthetically generated dataset demonstrate that the Hybrid solver yields statistically superior demand coverage compared to a SIC problem solved with the APOPT [^2] Advanced Process OPTimizer solver.
Also Wang et al. [157] solve the FLP. Their paper introduces the Variational Quantum Algorithm-Preserving Feasible Space (VQA-PFS) for solving the Uncapacitated Facility Location Problem (UFLP). VQA-PFS combines mixed operators and Hardware-Efficient Ansatz to enhance success probability and reduce circuit depth, outperforming QAOA, QAOA+, and HEA in efficiency and scalability.
A special case of FLP is the work by Chiscop et al. [33]. This paper presents a hybrid solution method for the Multi-Service Location Set Covering Problem (MSLSCP). It formulates the problem as a Quadratic Unconstrained Binary Optimization (QUBO) problem and uses D-Wave’s quantum annealer combined with classical optimization techniques to improve solution accuracy and performance.
A new popular topic here is the charging station location problem (CSLP). Radvand et al. [119] present a quantum search-based optimization algorithm for solving this problem. By utilizing Grover’s Adaptive Search and Quantum Phase Estimation, the algorithm offers a quadratic improvement in complexity over classical methods, optimizing charging station placement with enhanced efficiency in large transportation networks. Also Sakib et al. [127] provide a solution strategy for this problem using a QA approach. The model, validated through a real-world case study and solved on D-Wave quantum computers, demonstrates that QA can efficiently identify optimal EVCS locations, improving service quality for users.
@@
The paper by Klar et al. [76] addresses the challenges of factory layouts, emphasising their impact on operational costs. The problem involves assigning functional units to positions, optimising transportation distances. Manual layout planning is time-consuming due to the complexity of considering multiple parameters simultaneously. The study introduces a QA-based approach for factory layout planning, involving the formulation of the problem as a QUBO problem. The results show that QA can solve layout planning problems of different sizes within seconds, overcoming the trade-off between solution quality and computation time.